What is the value of the input Variable at which the maxima of function V = f(t) = IRsinωt occurs


1
Expert's answer
2022-02-02T11:45:04-0500

Let us find the value of the input variable at which the maximum of function "V = f(t) = IR\sinωt" occurs. It follows that "f'(t) = ωIR\cosωt=0" implies "ωt=\frac{\pi }2+\pi n", and thus "t=\frac{\pi }{2ω}+\frac{\pi n}{\omega}." Since

"f''(t) = -ω^2IR\sin ωt" and "f''(\frac{\pi }{2ω}+\frac{\pi n}{\omega}) = -ω^2IR\sin ω(\frac{\pi }{2ω}+\frac{\pi n}{\omega})
= -ω^2IR\sin (\frac{\pi }{2}+\pi n),"

we conclude that "f''(\frac{\pi }{2ω}+\frac{2\pi k}{\omega}) =
= -ω^2IR\sin (\frac{\pi }{2}+2\pi k)
= -ω^2IR\sin (\frac{\pi }{2})= -ω^2IR<0"

and "f''(\frac{\pi }{2ω}+\frac{\pi (2k+1)}{\omega}) =
= -ω^2IR\sin (\frac{\pi }{2}+\pi (2k+1))
= ω^2IR\sin (\frac{\pi }{2})= ω^2IR>0." Therefore, the points "t_k=\frac{\pi }{2ω}+\frac{2\pi k}{\omega}," where "k\in \N," is the values of the input variable at which the maximum of function "V = f(t) = IR\sinωt" occurs.